OpenStudy (anonymous):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2 h, and Car B traveled the distance in 1.5 h. Car B traveled 15 mph faster than Car A. How fast did Car B travel? mph

OpenStudy (anonymous):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2 h, and Car B traveled the distance in 1.5 h. Car B traveled 15 mph faster than Car A. How fast did Car B travel? : Let s = Car B's speed Car B is 15 mph faster than A, therefore (s-15) = Car A's speed : Write a distance equation; dist = speed * time 2(s-15) = 1.5s 2s - 30 = 1.5s 2s - 1.5s = 30 .5s = 30 s = 60 mph is car B's speed : : Confirm this find the distances, they should be equal 2(60-15)) = 90 1.5(60) = 90

OpenStudy (anonymous):

DANG thank you SO much

OpenStudy (anonymous):

Your welcome